OpenStudy (akiaki28):

consider an object with mass = 2kg moving at a velocity of v= 5m/s. It traverses a distance of 7m before coming to a halt due to friction. How strong is the frictional force acting on the object assuming that it is constant across the trajectory?

OpenStudy (osprey):

akiaki28 consider an object with mass = 2kg moving at a velocity of v= 5m/s. It traverses a distance of 7m before coming to a halt due to friction. How strong is the frictional force acting on the object assuming that it is constant across the trajectory? Kinematic equation for DECELERATION ... v squared = u squared - 2 a s 0=5 squared-2 a 7 25=2 a 7 a = 25/(2x7) The maths form of Newton's second law is often written F=ma. F = 2x25/(2x7) F = 25/7 F =3.5714285714285714285714285714286 I think/hope. http://perendis.webs.com

OpenStudy (shamim):

welcome to openstudy!!!

OpenStudy (osprey):

an kinematics and dynamics and Newton's laws and maybe ...

OpenStudy (shamim):

F=ma F=? m=mass is given u jv to find out a=?

OpenStudy (shamim):

now use v^2=u^2-2as

OpenStudy (shamim):

u,v, s. given a=?

likeabossssssss (likeabossssssss):

^^^^^

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)